# Do massive neutrino states really exist?

### v1

@@ 3.2 Neutrino Oscillation
Stray Forward to by deduce calculations witch will be reported elsewhere, considering processes with two flavor changes allows us to reproduce the standard formula for the probability of neutrino oscillations:
$P_{e\to\mu}\propto\cfrac{\sin^{2}(2\theta)}{\text{L}^{2}}\ \sin^{2}\left(\cfrac{\Delta m^{2}}{4E_{\text{avg}}}\text{L}\right),$ (3.7)
where values
$\Delta m^{2}=\text{tr}(\mathbb{M})\sqrt{\text{tr}^{2}(\mathbb{M})-4\det(\mathbb{M})}\quad\text{and}\quad\sin^{2}(2\theta)=\cfrac{2m_{e\mu}}{\sqrt{\text{tr}^{2}(\mathbb{M})-4\det(\mathbb{M})}}$ (3.8)
are expressed through the invariants of the mass matrix. The quantities $L$ are the distance between the source and the detector, and $E_{\text{avg}}$ are the average energies in the detector and the source.
@@ Conclusions
Consideration of processes with violation of lepton number showed that it does not matter whether we work in the mass or flavor states basis. We obtain the same expressions, while in weak interactions flavor states appear.
In this work, we derived the matrix propagator both in its general form and with explicit angular dependence. Also we write explicit mass matrix for 3 flavor through masses and angles. It appears to be qualitatively different in the cases of direct and inverse hierarchy, and the latter seems to exhibit a sort of fine tuning, which may be considered as an extra support in favor of direct hierarchy.
For the local diagram of the flavor-violating process $l_{\alpha}^{-}l_{\beta}^{+}\to W^{+}W^{-}$, we obtained a cross section of $\mathcal{O}(10^{-53})$, indicating no observable contribution to this process.
@@ Acknowledgments
Figure 4: Diagonal elements $m_{ee}$, $m_{\mu\mu}$, $m_{\tau\tau}$ of the neutrino mass matrix in NH versus $m_{1}$
@@
Figure 8: Diagonal elements $m_{ee}$, $m_{\mu\mu}$, $m_{\tau\tau}$ of the neutrino mass matrix in IH versus $m_{3}$

### v2

@@ 3.2 Neutrino Oscillations
Omitting detailed calculations, we find that considering processes with two flavor changes allows us to reproduce the standard formula for the probability of neutrino oscillations:
$P_{e\to\mu}\propto{\sin^{2}(2\theta)}\ \sin^{2}\left(\cfrac{\Delta m^{2}}{4E_{\text{avg}}}\text{L}\right),$ (3.12)
where the quantities
$\Delta m^{2}=\text{tr}(\mathbb{M})\sqrt{\text{tr}^{2}(\mathbb{M})-4\det(\mathbb{M})}\quad\text{and}\quad\sin(2\theta)=\cfrac{2m_{e\mu}}{\sqrt{\text{tr}^{2}(\mathbb{M})-4\det(\mathbb{M})}}$ (3.13)
are expressed through the invariants of the mass matrix.
@@ Conclusions
The consideration of processes with lepton number violation shows that it makes no difference whether one works in the mass or flavor basis: the same expressions are obtained, although flavor states appear in weak interactions.
Working in the flavor basis, we derived the matrix propagator both in its general form and with explicit angular dependence. This may seem like an overly elaborate construction for processes that can be easily diagonalized in the mass basis. However, in situations where diagonalization is not straightforward (for instance, in interactions with matter [5]), it becomes necessary to employ matrix propagators.
For the neutrino mass matrix, we obtained an explicit analytical expression in terms of the standard parameters: mixing angles, CP-violating phases, and neutrino masses. To our knowledge, such a representation has not been previously reported in the literature. Based on the global fit data from NuFIT 6.0 [3] and the joint T2K and NOvA analysis [4], we applied the Monte Carlo method to generate distributions of the mass matrix elements, thereby obtaining their uncertainties and numerical values for both normal and inverted hierarchies.
The data also allowed us to study the dependence of the mass matrix elements on the lightest neutrino mass for each hierarchy. Interestingly, in the inverted hierarchy, the real part of the element $m_{e\mu}$ is very small: its magnitude reaches a maximum at positive values and then asymptotically approaches zero. Given the large uncertainty in the estimate of this element, its precise value cannot be determined. Nevertheless, such behavior may indicate a fine-tuning that seems to flavor the normal hierarchy.
Next, we considered the crossing process of two oppositely charged leptons producing a pair of W bosons, $l_{\alpha}^{-}l_{\beta}^{+}\to W^{+}W^{-}$, within a two-flavor toy model. As expected, the contribution from neutrino exchange is extremely small, $d\sigma\simeq 10^{-11}\penalty 10000\ \text{fbarn}$. The cross-section for the electron-positron case $e^{+}e^{-}\xrightarrow{\nu}W^{+}W^{-}$ is seven orders of magnitude smaller than that reported in [6], since we did not include heavy neutrinos and considered only the neutrino-exchange contribution, which is suppressed relative to the dominant $e^{+}e^{-}\xrightarrow{Z,\ \gamma}W^{+}W^{-}$ annihilation into two W-bosons and their interference. It should be noted that the main contribution from the neutrino exchange process will be most apparent in the interference term with the annihilation process, as it depends linearly on this matrix element. This suggests searching for charge asymmetry in $W^{+}W^{-}$ final states, where interference between C-odd annihilation and neutrino-exchange amplitudes could produce observable effects linear in the weak amplitude, unlike the quadratically suppressed direct contribution.
The cross-section for the lepton flavor-violating process $e^{-}\mu^{+}\to W^{-}W^{+}$ demonstrates extreme suppression, reaching approximately $d\sigma\simeq 10^{-62}\penalty 10000\ \text{fb}$ - about 51 orders of magnitude smaller than comparable flavor-conserving processes. This pattern of extreme suppression aligns with findings in lepton number violating interactions; notably, [7] reported a difference of 59 orders of magnitude between Majorana and Dirac neutrino cross sections in similar processes. This dramatic discrepancy can be naturally explained within frameworks incorporating heavy neutrino states, particularly through the seesaw mechanism, where the inclusion of heavy Majorana neutrinos significantly enhances the amplitude of lepton number violating processes compared to the Dirac case.
The normal and inverted hierarchies exhibit opposite behaviors: in the normal hierarchy (NH), the cross-section increases with the lightest neutrino mass $m_{1}$, while in the inverted hierarchy (IH), it decreases with the lightest mass $m_{3}$. However, such differences cannot be tested experimentally due to the extremely small cross-sections.
Flavor-changing processes become apparent when neutrinos propagate over large distances, necessitating the consideration of macroscopic (long-distance) diagrams. The calculation of such a diagram within the QFT framework reproduces the well-known quantum mechanical formula for neutrino oscillations. Furthermore, the expressions obtained for the matrix propagator agree with those in [8], while in our case all formulas are expressed explicitly in terms of the invariants of the neutrino mass matrix.
@@ Acknowledgments
Figure 6: Diagonal elements $m_{e}$, $m_{\mu}$, $m_{\tau}$ of the neutrino mass matrix in NH versus $m_{1}$
@@
Figure 10: Diagonal elements $m_{e}$, $m_{\mu}$, $m_{\tau}$ of the neutrino mass matrix in IH versus $m_{3}$